OpenStudy (anonymous):

Simplify. 12 – 16 ÷ (–4) A. 1 B. 16 C. 8 D. –1

OpenStudy (idealist10):

12-16/-4 16/-4=-4 so 12-(-4)=12+4=16 Answer: 16

OpenStudy (anonymous):

thanks